### v2

# Field-induced states and thermodynamics of the frustrated Heisenberg antiferromagnet on a square lattice

Abstract
We investigate the ground-state and finite-temperature properties of the $J_{1}$-$J_{2}$ Heisenberg antiferromagnet on the square lattice in the presence of an external magnetic field. We focus on the highly frustrated regime around $J_{2}\approx J_{1}/2$. The $h$-$T$ phase diagram is investigated with particular emphasis on the finite-temperature transition into the “up-up-up-down” state that is stabilized by thermal and quantum fluctuations and manifests itself as a plateau at one half of the saturation magnetization in the quantum case. We also discuss the enhanced magnetocaloric effect associated to the ground-state degeneracy that arises at the saturation field for $J_{2}=J_{1}/2$. For reference, we first study the classical case by classical Monte Carlo simulations. Then we turn to the extreme quantum limit of spin-1/2 where we perform zero- and finite-temperature Lanczos calculations.
I Introduction
The Heisenberg antiferromagnet with nearest-neighbor, $J_{1}>0$, and next-nearest-neighbor bonds, $J_{2}\geq 0$, called $J_{1}$-$J_{2}$ model, or frustrated square lattice antiferromagnet (FSAFM), is the archetypical frustrated spin model [1, 2]. The corresponding Hamiltonian is given by
${\cal H}=J_{1}\sum_{\langle i,j\rangle}\bm{s}_{i}\cdot\bm{s}_{j}+J_{2}\sum_{\langle\!\langle i,j\rangle\!\rangle}\bm{s}_{i}\cdot\bm{s}_{j}-h\sum_{i}s^{z}_{i}$ (1)
including a coupling with an applied magnetic field $h$. Here, $\langle\cdot,\cdot\rangle$ denotes summation over the nearest-neighbor bonds $J_{1}$ whereas $\langle\!\langle\cdot,\cdot\rangle\!\rangle$ corresponds to the next-nearest neighbor $J_{2}$ bonds that couple spins on the diagonals of square plaquettes.
The $J_{1}$$J_{2}$ square-lattice model was introduced in the late 1980s with the aim to describe the breakdown upon doping of the Néel antiferromagnetic (NAF) state in the high-temperature cuprate superconductors [3, 4, 5]. Over the last few decades the zero-field case, $h=0$, has attracted much attention, since it is a canonical model to study frustration-induced quantum phase transitions between semiclassical ground-state phases with a magnetic long-range order (LRO) and disordered quantum phases, see, e.g., Refs. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59].
Based on numerous studies using a plethora of quantum many-body methods, consensus has been achieved that the $J_{1}$-$J_{2}$ model in the extreme quantum limit $s=1/2$ exhibits a magnetically disordered spin-rotation-invariant quantum phase in a small region $J_{2}^{c1}<J_{2}<J_{2}^{c2}$ around $J_{2}=0.5J_{1}$. The semi-classical ground-state phases of the model, namely the Néel antiferomagnetic LRO at small $J_{2}/J_{1}<J_{2}^{c1}$ and the collinear antiferromagnetic LRO at large $J_{2}/J_{1}>J_{2}^{c2}$, are well understood. Concerning the critical values $J_{2}^{c1}$ and $J_{2}^{c2}$, some estimates are $J_{2}^{c1}=0.44-0.46$ and $J_{2}^{c2}=0.59-0.6$ [20, 32, 36, 37, 40, 44, 47, 59]. However, some recent PEPS approaches [38, 42, 41, 52] lead to a transition point $J_{2}^{c1}$ close to $0.5$ or even above. Concerning the nature of the intermediate magnetically disordered quantum phase, there is an ongoing active controversial debate [17, 20, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 44, 45, 46, 48, 50, 52, 58]. A particular controversy concerns the very existence of an excitation gap in the intermediate quantum phase. A finite excitation gap was reported in Refs. [23, 27, 28, 33], whereas indications of a gapless spin liquid state were found in Refs. [29, 30, 36, 50]. Investigations based on density matrix renormalization group with explicit implementation of $SU(2)$ spin rotation symmetry in Ref. [32] report on a gapless spin liquid for $0.44<J_{2}/J_{1}<0.5$ and a gapped plaquette valence bond phase for $0.5<J_{2}/J_{1}<0.61$. This structure of the phase diagram is reaffirmed in Ref. [50], but the point separating the two intermediate phases is shifted to $J_{2}/J_{1}\approx 0.56$. On the other hand, a recent matrix-product state method claims that there is no spin-liquid phase in this model [52].
In addition to the basic theoretical interest in this nontrivial quantum many-body model, it also attracts attention thanks to its relation to experimental studies of various magnetic materials, such as VOMoO4 [60], $\mathrm{Li}_{2}\mathrm{VOSiO}_{4}$ and $\mathrm{Li}_{2}\mathrm{VOGeO}_{4}$ [61, 62, 63], BaCdVO(PO4)2 [64], RbMoOPO4Cl [65], SrLaCuSbO6 and SrLaCuNbO6 [66], NaZnVOPO4(HPO4) [67], YbBi2IO4 and YbBi2ClO4 [68], and KMMPO4 [69, 70]. However, to the best of our knowledge so far no magnetic material with an exchange-parameter ratio $J_{2}/J_{1}$ suitable for a magnetically disordered ground-state phase has has yet been found.
The unconventional quantum nature of the zero-field ground state in the strong-frustration regime has also stimulated the search for unconventional features in the magnetization curve [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83] as well as finite-temperature properties [16, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93] of the spin-1/2 FSAFM. In particular, in the regime of strong frustration $J_{2}/J_{1}\approx 0.5$, where the zero-field quantum ground state does not exhibit magnetic LRO, evidence for a well-pronounced plateau at $1/2$ of saturation was found [72].
In this paper, we revisit the behavior of the frustrated square-lattice antiferromagnet in an external magnetic field at zero and at finite temperature, where we focus on the regime around $J_{2}/J_{1}=0.5$. In Sec. II we first revisit the $h$-$T$ phase diagram of the classical Heisenberg model at $J_{2}/J_{1}=0.5$. Then, in Sec. III we switch to the extreme quantum limit of spin $s=1/2$. We first revisit the zero-temperature phase diagram in an external field in Subsec. III.1 and then study finite-temperature properties in an external field in Subsec. III.2. Finally, we conclude in Sec. IV with a summary and perspectives. We note that while much of the work on the spin-1/2 model focusses on the controversial $h=0$, $T=0$ phase diagram in the regime around $J_{2}/J_{1}=0.5$, here we do not intend to contribute to this particular question that is intensively studied by many other authors, as reviewed above.
II Classical model
II.1 Ground states
We begin with the classical $J_{1}$-$J_{2}$ model (1) treating spins $\bm{s}_{i}$ as three-component unit vectors. For the special ratio of the two exchange constants $J_{2}/J_{1}=1/2$, magnetic frustration is enhanced featuring an infinite classical degeneracy for all magnetic fields below the saturation [72]. In order to see this we define the total spin of a square plaquette $\alpha$:
$\bm{L}_{\alpha}=\sum_{i\in\alpha}\bm{s}_{i}\,$, (2)
as a sum over four surrounding sites. Then, for $J_{2}=J_{1}/2$, the spin Hamiltonian (1) can be rewritten as a function of $\bm{L}_{\alpha}$ variables only
${\cal H}=\frac{1}{4}\sum_{\alpha}\Bigl(J_{1}\bm{L}_{\alpha}^{2}-hL^{z}_{\alpha}\Bigr).$ (3)
So far, this is an exact transformation valid for any value of the spin. Let us now focus on the classical case. Minimizing energy of a single square plaquette we obtain from Eq. (3):
$L^{z}_{\alpha}=h/(2J_{1})\,.$ (4)
Any classical spin configuration that satisfies the above constraint for every square plaquette yields the lowest ground-state energy. The subsequent analysis of degenerate ground states splits into two parts: (i) finding equilibrium four-sublattice configurations for one block and (ii) extending single-block states to the entire lattice. Before doing that we note that the constraint (4) gives the following equilibrium magnetization $m=L^{z}_{\alpha}/4=h/(8J_{1})$, which in turn yields for the saturation field $h_{\rm sat}=8J_{1}$. At $T=0$, the classical magnetization curve $m(h)$ is a straight line up to $h_{\rm sat}$.
Figure 1: Spin configurations minimizing the classical energy of a single four-site plaquette of the FSAFM in different magnetic fields. The collinear uuuduuud state produces the 1/2 magnetization plateau.
The constraint (4) imposes three conditions on the eight angle variables for the four spins of each block. One more angle variable is eliminated by the broken $O(2)$ rotation symmetry about the field direction. Overall, this counting yields four continuous variables to describe various classical ground states for a single block. Analysis of the effect of fluctuations, thermal or quantum, on a state selection can be performed using the concept of an effective biquadratic exchange, see, for example, Ref. [94]. Figure 1 sketches the relevant spin configurations that appear for four-site blocks in various magnetic fields [72, 95, 96]. The “up-up-up-down” (uuuduuud) configuration has the lowest energy for magnetic fields around $h_{\rm sat}/2$ and produces the magnetization plateau with $m=1/2$. At higher fields away from the 1/2 plateau, spins tilt retaining the 3:1 splitting. For fields below $m<1/2$, the noncollinear spin configuration has a more complex 2:1:1 structure. The 2:2 state, which is one of the classical ground states in the entire range $0<h<h_{\rm sat}$, is stable, in the presence of a biquadratic exchange, in the low-field region only [96].
Since the plaquettes $\alpha$ share edges, one spin participates in several plaquettes. As a result, one faces the problem of extending a single-block solution to the entire lattice. This leads to additional degeneracy for the lattice spin model for each of the single-block solutions in Fig. 1. The corresponding degeneracy is most easily visualized for the uuuduuud state [72]. Specifically, consider the most symmetric uuuduuud configuration. For that we choose the down spin, for example, in the lowest left corner of the first square plaquette and extend this configuration in a translationally invariant way over the whole lattice. This state consists of parallel, vertical and horizontal, lines of up-down spins separated by lines of up spins. The alternating up-down spins along any of these lines can be freely rotated around an arbitrary angle without changing the total classical energy [72]. Since the uuuduuud state permits a maximum of such deformations, one may expect it to be selected by thermal fluctuations via the “order-by-disorder” mechanism [97, 98, 99, 100, 101, 102]. A subtle property of the order-by-disorder mechanism is a competition between a local fluctuation contribution manifested by the effective biquadratic coupling and the extended zero-energy modes related to spin distribution over the entire lattice. Furthermore, in the latter case there is the possibility of distinct effects from the thermal fluctuations, which rely on the entropic contribution of the low-energy (zero) modes, and from the quantum fluctuations, with magnons at different momenta contributing equally. As we shall see below, all these exotic features of the order-by-disorder effect are relevant for the FSAFM.
Figure 2: Field-temperature phase diagram of the classical FSAFM with $J_{2}=0.5$, $J_{1}=1$. Magnetic states with (quasi) LRO are labeled according to Fig. 1, PM stands for a disordered paramagnetic state. Phase transition points obtained from the Monte Carlo simulations are shown by open circles with solid lines as a guide for the eye. The dashed lines indicate the expected Berezinskiĭ-Kosterlitz-Thoulesss transition boundaries.
II.2 Field-temperature phase diagram
Next, we present new classical Monte-Carlo results on significantly larger lattices compared to the previous studies [72, 74]. Simulations have been performed using the single-spin flip Metropolis algorithm in combination with the microcanonical over-relaxation moves. Details of the algorithm are provided in recent studies devoted to various classical frustrated spin models [103, 104, 105].
Figure 2 presents the constructed $h$-$T$ phase diagram for $J_{2}/J_{1}=1/2$. First, we observe the expected stabilization of the uuuduuud phase by thermal fluctuations in the regime $T\lesssim J_{1}/10$. The stability region is asymmetric around $h_{c}=0.5h_{\rm sat}$ having a significant extent into the low-field part of the diagram. Second, there are sizable regions with the 2:1:1 and 3:1 states just below or above the uuuduuud phase, respectively. The two noncollinear states possess the longitudinal LRO in the $s^{z}$ components accompanied by the quasi-LRO in transverse spin components. Mapping this from magnetic to bosonic language [106, 107, 108], these two phases can be identified as supersolids, whereas the parent uuuduuud phase is an insulating state with the density-wave crystal.
Figure 3: The magnetization curve $m(h)$ (lines, right axis) and field derivative of the magnetization (symbols connected by lines, left axis) for the classical FSAFM with $J_{2}=0.5$, $J_{1}=1$ at three different temperatures obtained on lattices with $L=32$.
Figures 3 and 4 illustrate some of the data used in the construction of the phase diagram Fig. 2. According to the analysis at the beginning of this section, the magnetization curve $m(h)$ at zero temperature is a straight line until the saturation field $h_{\rm sat}$. Figure 3 shows some magnetization curves at low but finite temperature. The main feature that emerges at intermediate fields is a small dip for $m\lesssim 1/2$ that is most pronounced for $T/J_{1}=0.03$. We note that the improved Monte Carlo results do not reproduce the hysteresis that appeared in earlier work [72]. To better exhibit these features in the magnetization curve, we consider the field derivative of the magnetization ${\rm d}m/{\rm d}h$, also shown in Fig. 3. At zero temperature $m(h)$ has a constant slope for all fields below the saturation field $h_{\rm sat}$, i.e., ${\rm d}m/{\rm d}h=J_{1}/8$ at $T=0$ and for $h<h_{\rm sat}$. At $T>0$, we observe in Fig. 3 the emergence of small but sharp and distinct maxima in ${\rm d}m/{\rm d}h$, which may be taken as first indications of field-induced phase transitions. For $h\lesssim 4\,J_{1}=h_{\rm sat}/2$, we observe a drop in ${\rm d}m/{\rm d}h$. The slope of the magnetization curve $m(h)$ remains finite and thus one finds only a small dip but no strict plateau in the classical magnetization curve. Nevertheless, the reduced slope of the magnetization curve can be considered as a distinctive feature of the uuuduuud phase.
[…]
To further characterize different phases, we look now at the static spin structure factor $S^{\alpha\beta}(\bm{Q})$:
$S^{\alpha\beta}(\bm{Q})=\frac{1}{N^{2}}\>\sum_{i,j}\langle s^{\alpha}_{i}s^{\beta}_{j}\rangle\,e^{i\bm{Q}\cdot(\bm{R}_{i}-\bm{R}_{j})}\,$, (5)
[…]
Figure 5: Field derivative of magnetization for the classical FSAFM at $T=0.06$ for different exchange ratios close to $J_{2}=0.5$ (in units of $J_{1}=1$).
Figures 5 and 6 illustrate stability of the uuuduuud state upon moving away from the strongly frustrated point $J_{2}\neq 0.5J_{1}$. Focusing first on field scans in Fig. 5 for $T=0.06J_{1}$, one can clearly observe sharp anomalies in the field derivative of the magnetization ${\rm d}m/{\rm d}h$ for $J_{2}/J_{1}=0.5\pm 0.005$ indicating the upper and lower boundaries of the the uuuduuud plateau state. In contrast, ${\rm d}m/{\rm d}h$ remains completely flat for $J_{2}/J_{1}=0.51$ featuring canted Néel and stripe states in the entire range of magnetic fields. For $T=0.06J_{1}$ and $h=3.6J_{1}$, Fig. 6 shows the dependence of the transverse and longitudinal components of the static structure factor upon varying the ratio of $J_{2}/J_{1}$. Here, one can see the same trend with the uuuduuud state clearly stabilized by thermal fluctuations in a finite range of $J_{2}/J_{1}$. Note that the uuuduuud state is the classical ground state only for a single frustration ratio $J_{2}/J_{1}=1/2$. Hence, for $J_{2}/J_{1}\neq 1/2$, the region of the uuuduuud phase is detached from the vertical axis of the phase diagram at $T=0$ and requires sufficiently strong thermal fluctuations for its stabilization. Conversely, for $J_{2}/J_{1}=1/2$ the strong thermal fluctuations for $T\gtrsim 0.1J_{1}$ destroy the uuuduuud LRO, see Fig. 2. Thus, one requires a delicate balance of sufficiently but not too strong thermal fluctuations to stabilize the uuuduuud state, consistent with its disappearance when $J_{2}/J_{1}$ deviates by about $0.01$ from the degenerate ratio $1/2$.
Figure 6: Static structure factor of the classical FSAFM for fixed $T=0.06$ and $h=3.6$ (in units of $J_{1}=1$) as a function of $J_{2}/J_{1}$ for three lattice sizes $L$.
III Quantum model
We start with a basic consideration that is valid for any spin quantum number $s$: the single-magnon dispersion above the ferromagnetically polarized background is a single-particle problem that is straightforward to solve. One finds the single-magnon dispersion to be given by (compare also Ref. [73])
$\omega(k_{x},k_{y})$ $=$ $2s\bigl(J_{1}\left(\cos{k_{x}}+\cos{k_{y}}\right)$ (6)
$\qquad+J_{2}\left(\cos(k_{x}+k_{y})+\cos(k_{x}-k_{y})\right)\bigr)$
$-4s\left(J_{1}+J_{2}\right)+h$
$=$ $2s\left(J_{1}\left(\cos{k_{x}}+\cos{k_{y}}\right)+2J_{2}\cos{k_{x}}\cos{k_{y}}\right)$
$-4s\left(J_{1}+J_{2}\right)+h\,.$
If the transition to saturation is a second-order quantum phase transition, the minimal single-magnon energy vanishes exactly at the saturation field $\omega(\bm{k}_{\min})-h_{\rm sat}=0$. Applying this condition to Eq. (6), we can determine the saturation field as
$h_{\rm sat}=\begin{cases}8sJ_{1}&\text{for }J_{2}\leq J_{1}/2,\\ 4s(J_{1}+2J_{2})&\text{for }J_{2}\geq J_{1}/2.\end{cases}$ (7)
In particular $h_{\rm sat}=8sJ_{1}$ at $J_{1}=J_{2}/2$. In the classical limit $s\to\infty$ and upon suitably rescaling the spins by $1/s$, i.e., by effectively replacing $s\to 1$, we find $h_{\rm sat}=8J_{1}$ at $J_{1}=J_{2}/2$ in the classical limit, as already stated in Subsec. II.1.
An interesting special case is $J_{2}=J_{1}/2$ where the dispersion relation (6) simplifies (see also Ref. [77])
$\left.\omega(k_{x},k_{y})\right|_{J_{2}=J_{1}/2}$ $=$ $2sJ_{1}\left(\left(1+\cos k_{x}\right)\left(1+\cos k_{y}\right)-4\right)$ (8)
$+h$
such that
$\left.\omega(k_{x},\pi)\right|_{J_{2}=J_{1}/2}=\left.\omega(\pi,k_{y})\right|_{J_{2}=J_{1}/2}=-8sJ_{1}+h\,$, (9)
i.e., we find degenerate lines of minima for $h\geq h_{\rm sat}$. This is reminiscent of the classical degeneracy discussed in Subsec. II.1, in particular the degenerate manifold evolving from the uuuduuud state at $m=1/2$.
Figure 7: The $J_{1}$-$J_{2}$ model on a finite square lattice of $N=52$ sites. Repeated sites (shown in blue) indicate periodic boundary conditions.
III.1 Zero-temperature $s=1/2$ phase diagram
To numerically investigate the quantum case at finite magnetization $m<1$, we now restrict to the extreme quantum limit $s=1/2$. We employ the Lanczos method [112, 113]
Figure 8: Zero-temperature magnetization curves $m(h)$ of the spin-1/2 FSAFM for $J_{2}/J_{1}=0$ (a), $0.5$ (b), $0.55$ (c), and $0.6$ (d). Full curves denote finite-size results for the given system size $N$. The dashed red lines indicate an estimate for the behavior of the infinite system. This estimate is obtained by connecting the midpoints of the finite-size steps for the largest available system size NN [114], except for suspected plateaux (including the transition to saturation $m=1$), where we use the corners of the steps instead. For a discussion why we draw $m=1/2$ plateaux in the extrapolated curves at $J_{2}/J_{1}=0.55$ and $0.6$, but none for $J_{2}/J_{1}=0.5$, compare Fig. 9 below and the discussion in the text; for the identification of this $m=1/2$ plateau with the uuuduuud state, see Fig. 2 below and the discussion in the text.
to find the minimal eigenvalues and -vectors of the Hamiltonian and thus determine its ground-state properties. We use in particular $S^{z}$ conservation. The total Hilbert space for a given $S^{z}$ and $N$ spins 1/2 has dimension $N\choose N/2-S^{z}$}. Consequently, the dimension is maximal for $S^{z}=0$ and decreases with increasing $S^{z}$, i.e., higher magnetic fields. In particular, we can push computations to the $N=52$ lattice shown in Fig. 7 for $m\gtrsim 1/2$, i.e., $S^{z}\gtrsim N/4=13$. This covers in particular the region relevant to a possible $m=1/2$ plateau. Note that this $N=52$ lattice shares the $90^{\circ}$ rotational symmetry of the infinite lattice, but it breaks spatial reflection symmetries.
III.1.1 Magnetization curves
Figure 8 presents zero-temperature magnetization curves, where we normalized the saturation magnetization to $m=1$. In addition to finite-size results, we show estimates for the behavior of the infinite system by the dashed red lines. Following Ref. [114], this estimate is in general obtained by connecting the midpoints of the finite-size steps for the largest available system size $N$. The exception are suspected plateaux: here we use the corners of the steps instead. The main bias in this extrapolated curves is whether or not we expect a plateau and we refer to the discussion below for further details on this point.
For reference, we include the plain square-lattice case with $J_{2}=0$ (Fig. 8(a)) that has been investigated in detail previously, see for example Ref. [115], including the $N=40$ results. Here, we find a smooth curve in the thermodynamic limit with an upwards curvature that ends in a logarithmic singularity. This behavior is characteristic of a two-dimensional quantum spin system, and has been analytically predicted for the present case by second-order spin-wave theory [116].
Figure 8(b) presents the results for the highly frustrated case $J_{2}/J_{1}=1/2$. Here, we have no evidence for particular features around $m=1/2$ and thus have drawn a smooth curve as estimate for the behavior in the limit $N\to\infty$. However, the transition to saturation is exceptionally steep in this case, as is demonstrated by the inset to Fig. 8(b). This behavior can be traced to the degeneracy of the one-magnon spectrum arising at $J_{2}/J_{1}=1/2$, see Eq. (9). This degeneracy gives rise to excitations localized along horizontal or vertical lines across the lattice [75] and thus leads to a finite-size jump to saturation of size $\delta m=1/L$ for an $L\times L$ square geometry. In the thermodynamic limit, this is expected to give rise to a square-root singularity with a logarithmic correction at the saturation field $h_{\rm sat}$ [77].
Next, we consider the case $J_{2}/J_{1}=0.55$, see Fig. 8(c). Here, the $N=48$ system appears to be an outlier, but the other system sizes converge to a stable $m=1/2$ plateau. The resulting magnetization curve is in good agreement with results from the density matrix renormalization group (DMRG) method [81], and even the estimates of the lower and upper critical fields of the $m=1/2$ plateau are in reasonable quantitative agreement. Note furthermore that the slope of the magnetization curve is very steep when it approaches $m=1/2$ from below. This might be a signature of a first-order quantum phase transition, although the available data does not suffice to draw a definite conclusion on the nature of this field-induced phase transition.
Finally, we look at $J_{2}/J_{1}=0.6$. The resulting magnetization curve, shown in Fig. 8(d), is qualitatively similar to the case $J_{2}/J_{1}=0.55$. In particular, we find again an $m=1/2$ plateau that might be even broader than in the previous case. Our results are again consistent with the DMRG ones [81] for $J_{2}/J_{1}=0.6$.
We note that a Chern-Simons theory predicted a distinct $m=1/3$ plateau in addition to the $m=1/2$ one [76]. The $N=36$ and $48$ systems indeed exhibit a broader step for $J_{2}/J_{1}=0.55$ (Fig. 8(c)) and $0.6$ (Fig. 8(d)) while the magnetization curves for the other system sizes appear to exhibit a smooth behavior in this region. The appearance of a strict $m=1/3$ plateau is in fact only possible for system sizes divisible by three. It is difficult to say on the basis of the present data if sizes $N$ that are multiplies of three favor a finite-size plateau or if sizes not commensurate with three miss a one-third plateau. We therefore chose to follow Ref. [81] and have not drawn an $m=1/3$ plateau in our extrapolated curves.
Lastly, we note that at $J_{2}/J_{1}=0.5$ and $J_{2}/J_{1}=0.55$ we have drawn an $m=0$ plateau in Fig. 8(b) and (c), respectively. This corresponds to a spin gap in zero field, as is to be expected for the paramagnetic phases around $J_{2}/J_{1}=1/2$ mentioned in the Introduction. The situation for $J_{2}/J_{1}=0.6$ (Fig. 8(d)) is less clear. Here, the $N=36$ system suggests the presence of a spin gap, but this is not confirmed by the $N=40$ system. We therefore have not drawn an $m=0$ plateau although we cannot exclude the existence of a small $m=0$ plateau, i.e., spin gap for $J_{2}/J_{1}=0.6$. In any case, we recall that the zero-field behavior is a challenging issue that is beyond the scope of the present work.
Figure 9: Width $\Delta h$ of the $m=1/2$ step in the magnetization curve as a function of the frustration parameter $J_{2}/J_{1}$. Symbols represent actual data points whereas lines are guides to the eye.
Figure 10: Real-space correlations on the $N=52$ lattice for $m=1/2$ ($S^{z}=13$) as a function of $J_{2}/J_{1}$. Panel (a) shows the connected logitudinal correlations, panel (b) the transverse ones [^2] Reliable information about the corresponding values of $R$ is no longer available, as neither is the relevant author. We therefore refrain from labeling the curves and mention just that the same symbols and colors are used for transverse and longitudinal correlations at the same $R$. . Symbols represent actual data points whereas lines are guides to the eye.
III.1.2 $m=1/2$ plateau
Now we focus on a more detailed characterization of the $m=1/2$ plateau. First, we look at the width of the $m=1/2$ step, calculated as $\Delta h=E(N/2+1)-2E(N/2)+E(N/2-1)$, where $E(S^{z})$ is the energy of the ground state for the given value of $S^{z}$ at $h=0$. Figure 9 presents the behavior of $\Delta h$ as a function of $J_{2}/J_{1}$ for systems with $32\leq N\leq 52$ spins 1/2. We have chosen lattices with an aspect ratio as close as possible to unity for a given size $N$, but nevertheless the results fall into two groups, namely the square respectively rectangular geometries for $N=36$ and $48$ on the one side, and the tilted cases $N=32$, $40$, and $52$ on the other side. This non-monotonic finite-size behavior renders an extrapolation to $N\to\infty$ difficult. Nevertheless, we observe a clear and consistent enhancement of the $m=1/2$ step size in the region $0.52\lesssim J_{2}/J_{1}\lesssim 0.7$, which we thus take as a first estimate for the stability region of the $m=1/2$ plateau.
To gain further insight into the state of the $m=1/2$ plateau, Fig. 2 shows real-space spin-spin correlations at distance $R$ for the $N=52$ lattice of Fig. 7, i.e., at $S^{z}=13$. Note that $\langle s^{z}_{R}\rangle=m/2=1/4$. We therefore subtract a background of $\langle s^{z}_{0}\rangle\langle s^{z}_{R}\rangle=1/16$ from the longitudinal correlation function $\langle s^{z}_{0}s^{z}_{R}\rangle$ to obtain the connected one, that is shown in Fig. 2(a). The transverse one in Fig. 2(b) is obtained as $\langle s^{\perp}_{0}s^{\perp}_{R}\rangle=\langle\bm{s}_{0}\cdot\bm{s}_{R}\rangle-\langle s^{z}_{0}s^{z}_{R}\rangle$. Overall, we observe three distinct regimes in Fig. 2: (i) For $J_{2}/J_{1}\lesssim 0.5$ the transverse correlations dominate. This is consistent with Néel order in the transverse components, as expected in the limit $J_{2}=0$. (ii) For $0.52\lesssim J_{2}/J_{1}<0.7$, the transverse correlations essentially vanish, but strong longitudinal ones develop. This is consistent with the collinear uuuduuud state that is expected for an $m=1/2$ plateau from the investigation of the classical model in Sec. II. (iii) For $J_{2}/J_{1}>0.7$, the longitudinal components vanish again while a new signal develops in the transverse components. Different signs reflect an order that differs from the one in regime (i). Indeed, for $J_{1}=0$, one has two interpenetrating square lattices, each of which develops independent Néel order. A weak coupling $J_{1}>0$ then leads to a “stripe” phase, consistent with the correlations observed in Fig. 2(b) for $J_{2}/J_{1}>0.7$.
Taken together, the data for $\Delta h$ and the spin-spin correlation functions at $m=1/2$ are consistent with an uuuduuud plateau in the regime $0.52\lesssim J_{2}/J_{1}\lesssim 0.7$, gapless transverse Néel order for $J_{2}/J_{1}\lesssim 0.52$, and gapless transverse stripe order for $J_{2}/J_{1}\gtrsim 0.7$. These results on the stability of the uuuduuud state for the $s=1/2$ model are in sharp contrast with the thermal order by disorder for classical spins obtained above in Sec. II. Specifically, for a fixed $T$ the uuuduuud state appears rather symmetrically for $0.492\lesssim J_{2}/J_{1}\lesssim 0.51$ in the classical model, see Fig. 6. In fact, this behavior illustrates a remarkable distinction between quantum and thermal order-by-disorder effects already mentioned in the end of Sec. II.1. Using a so-called augmented spin-wave theory for large spins, Coletta et al. [80] have demonstrated such an asymmetric stabilization of the 1/2 plateau for $J_{2}/J_{1}\gtrsim 0.5$. This effect is driven by the high-energy magnons with $\omega(\bm{k})\approx Js$, which contribute to the vacuum zero-point energy on the same footing with the low-energy magnons. On the other hand, the thermal order by disorder is completely dominated by the low-energy magnons with $\omega(\bm{k})\approx T$.
III.2 Thermodynamics for $s=1/2$
Finally, we investigate the finite-temperature behavior of the $s=1/2$ FSAFM. This is motivated by two questions. Firstly, we are interested in the stability of the uuuduuud state at finite temperature and secondly we intend to verify the impact of the competing interactions on the magnetocaloric effect, in particular around the degeneracy arising at the saturation field $h_{\rm sat}$ for $J_{2}/J_{1}=1/2$. To address these questions, we focus on the two cases $J_{2}/J_{1}=0.5$ and $0.55$.
In order to obtain results for system sizes $N$ that are comparable to those that we have investigated at $T=0$, we employ the finite-temperature Lanczos method [117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127]. To be precise, we have employed two different implementations of this method. A first one follows the historic route [117, 118, 120, 121, 122, 123, 124] whereas a second one builds on the notion of “thermal pure quantum states” (TPQ) [128, 129]. With the Krylov approximation to the imaginary-time evolution [130], the two methods become equivalent [131], but the technical implementations are still different, and the second point of view lends a more solid theoretical foundation to the earlier heuristic approach. In order to distinguish the two methods, we refer to the historical approach as “FTLM” and to the more recent implementation [125] as “TPQ”.
Figure 11: Specific heat per spin of the $s=1/2$ FSAFM at $J_{2}/J_{1}=0.55$ for $h/J_{1}=4$ (a) and $h/J_{1}=2.4$ (b). Panel (a) compares three different methods for the $N=36$ system. Low-temperature features likely reflect the finite size of the system in this case. Panel (b) presents results for four different system sizes. In this case, features for $T\lesssim 0.02J_{1}$ are likely artifacts of either the finite size or the computational method whereas the peak for $T\approx 0.1J_{1}$ is taken as a signature of the phase transition into the $m=1/2$ uuuduuud state.
III.2.1 Specific heat
Figure 11 presents the specific heat per spin $C$ as a function of temperature at $J_{2}/J_{1}=0.55$ for two selected fields $h/J_{1}=4$ and $2.4$. TPQ and FTLM are subject to statistical errors that arise from sampling of the start vectors, corresponding to the $T=\infty$ limit. The cyan error tube around the $N=36$ TPQ result has been obtained with a jackknife analysis [132], whereas no error bars are shown for the FTLM results in Fig. 11.
The first case $h/J_{1}=4$ lies just below the saturation field that is $h_{\rm sat}/J_{1}=4.2$ for $J_{2}/J_{1}=0.55$ (see Eq. (7)). The result for the $N=36$ system is shown in Fig. 11(a) and serves as a benchmark. The proximity to saturation permits us to perform exact diagonalization (ED) that covers the complete spectrum up to sufficiently high energies to yield essentially exact results for the entire temperature range shown. This ED curve is free from the statistical errors that are inherent to FTLM and thus serves as a reference. We see that the FTLM curve is very close to the ED one, and also the TPQ curve deviates from these two only by an amount that is consistent with the estimated error. The features observed in Fig. 11(a) around $T/J_{1}=0.01$ and $0.05$ are most likely due to finite-size effects reflecting the finite system size $N=36$ and thus have no physical meaning.
Figure 11(b) shows results for $h/J_{1}=2.4$. This value of the field is chosen to lie in the middle of the $m=1/2$ plateau of the $T=0$ magnetization curve, see Fig. 8(c). TPQ and FTLM for the $N=36$ system are again consistent with each other. Beyond this, here we include results for other system sizes, namely $N=24$ ED and also FTLM for $N=40$ and $48$. We note that for $N\leq 40$ we have sampled all $S^{z}$ sectors, but for $N=48$ this is no longer possible. Validity of the $N=48$ FTLM results is thus restricted to high magnetic fields and low temperatures. In particular for $h/J_{1}=2.4$, we have to restrict the $N=48$ FTLM curve to $T/J_{1}\leq 0.15$. Nevertheless, all curves in Fig. 11(b) exhibit a maximum in $C(T)$ around $T/J_{1}\approx 0.1$. The system sizes studied here are larger than those of previous investigations [84, 85, 86, 87, 88, 89, 91], but they are still too small to permit an unambiguous determination of a thermodynamic phase transition. Nevertheless, the uuuduuud state breaks translational symmetry such that we expect a finite-temperature phase transition out of it, as in the classical version of the model. The maxima of the specific heat in Fig. 11(b) may serve as an estimate for the location of the phase transition such that we infer a transition temperature of the order of $T=J_{1}/10$ for $h/J_{1}=2.4$. By contrast, the upturns observed at low temperature in the $N=40$ and $48$ FTLM curves are likely to be artifacts of the method, while the true low-temperature symptotics of $C(T)$ is expected to be an exponential decrease for $T\to 0$ on any system of a finite size $N$.
Figure 12: Specific heat per spin $C$ (a,b) and entropy per spin $S$ (c,d) as a function of temperature and magnetic field for $J_{2}/J_{1}=0.5$ (a,c) and $0.55$ (b,d). These figures are composites of different system sizes, starting with $N=40$ for $m\lesssim 1/2$, going over $N=48$ for $1/2\lesssim m\lesssim 1$, $N=64$ for $m\gtrsim 1$ at high $T$ to $N=400$ for $m\gtrsim 1$ at low $T$. The white solid lines in panels (c,d) trace values of constant entropy per spin. The leftmost curve is for $S=0.05$, and $S$ increases in steps of $0.05$ with increasing temperature.
Finally, in Fig. 12 we show temperature-field scans of the specific heat per spin $C$ (upper row of panels) and the entropy per spin $S$ (lower row of panels) for $J_{2}/J_{1}=0.5$ (left column of panels) and $J_{2}/J_{1}=0.55$ (right column of panels). Here we have exploited the fact that the Hilbert space dimension decreases with increasing $m$ to build a composite of different system sizes such that we start with $N=40$ at low fields and increase $N$ with increasing $m$ and decreasing temperatures such that we reach $N=400$ for $m\gtrsim 1$ at low temperatures.
According to Subec. III.1.2, there should be no $m=1/2$ plateau at $T=0$ for $J_{2}/J_{1}=0.5$. If the case $s=1/2$ is analogous to the classical case shown in Fig. 2, the maximum of $C$ at $T/J_{1}\approx 0.1$ for $1\lesssim h/J_{1}\lesssim 2$ observed in Fig. 12(a) might nevertheless indicate a transition out of an uuuduuud phase stabilized by thermal fluctuations. For $J_{2}/J_{1}=0.55$, we find an $m=1/2$ plateau in the range of fields $2.24\lesssim h/J_{1}\lesssim 2.60$, see Fig. 8(c). The maximum of $C$ for $T/J_{1}\lesssim 0.1$ that one observes in this range of fields in Fig. 12(b) probably corresponds to a transition out of this plateau state, as we have discussed in the context of Fig. 11(b). Invoking again the analogy with the classical case Fig. 2, the continuation of this maximum to lower fields $1.2\lesssim h/J_{1}\lesssim 2.2$ and at $T/J_{1}\lesssim 0.08$ might again correspond to a transition out of an uuuduuud state stabilized by thermal fluctuations. This suggests an intriguing scenario where the crossover from stabilization by quantum to thermal fluctuations around $h/J_{1}\approx 0.55$ would be accompanied by a minimum in the transition temperature.
In the low-field limit, the specific heat of Fig. 12(a,b) exhibits a maximum around $T/J_{1}\approx 0.2$. This could be just a fluctuation maximum. However, it could also be a signature of a finite-temperature phase transition for a plaquette or columnar valence-bond solid ground state as suggested, e.g., in Refs. [33, 41, 52, 55, 59]. Such a state would break lattice symmetries (translations and/or rotations) and could thus support a finite-temperature phase transition. However, the features observed at $h=0$ in Fig. 12(a,b) are broad. If they correspond to phase transitions, finite-size effects would thus be large. In any case, clarification of the exact meaning of these low-field low-temperature maxima in the specific heat is a question that goes beyond the scope of the present work.
III.2.2 Entropy and magnetocaloric effect
We now turn to the entropy $S$ that is shown in Fig. 12(c,d). This quantity and the related magnetocaloric effect were investigated previously in the classical case [89] and for spin 1/2 [86, 87, 88, 91]. However, these previous investigations of the spin-1/2 model were restricted to $N\leq 24$ and consequently to relatively high temperatures. Since the entropy is related to the specific heat $C/T$ by an integral, the information content encoded in the entropy is equivalent to the specific heat that we discussed previously. In this respect, the $S=0.05$ isentrope, i.e., the leftmost white curve in Fig. 12(c,d) can be considered to roughly trace the expected transition into the uuuduuud phase that we have discussed above. However, the entropy also yields a more direct access to the magnetocaloric effect, as the isentropes show the behavior of the system under adiabatic changes of the magnetic field. We observe distinct field-controlled cooling around the saturation field in Fig. 12(c,d) and close to the quantum-phase transitions [133, 134] terminating the $m=1/2$ plateau in Fig. 12(d).
Geometric frustration is expected to enhance the magnetocaloric effect [135, 136, 137, 134]. In the present case, we expect such an enhancement specifically at $J_{2}/J_{1}=1/2$ where the single-magnon spectrum becomes degenerate, recall Eq. (9). Such an enhancement was indeed observed in previous works [86, 87, 91] that scanned the ratio $J_{2}/J_{1}$ at a fixed temperature. Here, such an enhancement is observed as a pronounced minimum in the lowest three isentropes $T(h)$ around $h_{\rm sat}$ in Fig. 12(c).
IV Conclusion
We have revisited the behavior of the $J_{1}$-$J_{2}$ FSAFM in a magnetic field, pushing previous investigations [72, 73, 74] to larger system sizes and higher accuracy.
We first determined the phase diagram of the classical Heisenberg model at $J_{2}/J_{1}=1/2$. Here, we confirmed order-by-disorder stabilization of the uuuduuud state around half of the saturation field and for temperatures $0<T\lesssim J_{1}/10$. In addition, we showed that 2:1:1 and 3:1 supersolid phases emerge below and above the uuuduuud one, respectively. Furthermore, we provided evidence that the uuuduuud phase persists in a small window around $J_{2}/J_{1}=1/2$, although the full phase diagram of the classical FSAFM remains to be investigated for $J_{2}/J_{1}\neq 1/2$.
Then we turned to the spin-1/2 case. First, we presented zero-temperature magnetization curves for several ratios of $J_{2}/J_{1}$ and provided evidence for the emergence of a distinct $m=1/2$ plateau that can be understood as an uuuduuud state stabilized by quantum fluctuations. We find a relatively large stabilization regime $0.52\lesssim J_{2}/J_{1}\lesssim 0.7$ that is shifted to values $J_{2}>J_{1}/2$ such that analytic treatment of this uuuduuud $m=1/2$ plateau requires expansion around a classically unstable state [80]. Intriguingly, the uuuduuud state is found to be stabilized by quantum fluctuations in the regime $h\gtrsim h_{\rm sat}/2$, whereas the stabilization by thermal fluctuations takes place at lower fields $h\lesssim h_{\rm sat}/2$. Such a trend is also known for the Heisenberg antiferromagnet on the triangular lattice, see, e.g., discussion in Ref. [104]. We expect that the supersolid phases found in the classical case below and above the uuuduuud phase exist also for $s=1/2$. While this expectation has been formulated earlier [74], a detailed investigation goes beyond the scope of the present work. Nevertheless, we mention that Ref. [81] found several non-trivial ordered states in the magnetization curves of the $s=1/2$ FSAFM, although it is not entirely clear to us if their Y-like and V-like phases match our 2:1:1 and 3:1 states below and above the plateau, respectively. Another point that might merit further attention is the question of the existence of the $m=1/3$ plateau predicted by a Chern-Simons theory [76]. A systematic numerical investigation would require a sequence of system sizes divisible by three, but here we have only two of those at our disposal, namely $N=36$ and $48$.
Lastly, we computed finite-temperature properties of the $s=1/2$ FSAFM in a magnetic field, relying mainly on FTLM. We found preliminary signatures in the specific heat of a finite-temperature phase transition into the uuuduuud state. Like in the classical case, these finite-temperature ordering transitions remain restricted to low temperatures $T/J_{1}\lesssim 0.1$. We further observed that the degeneracy in the single-magnon spectrum at $J_{2}/J_{1}=1/2$ leads to an enhanced magnetocaloric effect at the saturation field.
Acknowledgements.
Johannes Richter was a prolific scientist. Regrettably, he did not witness completion of all of his projects. This is one of those that remained unfinished. Johannes Richter contributed the zero-temperature Lanczos calculations, part of the finite-temperature Lanczos results, and participated in the early stages of preparing the manuscript. We are grateful to Jörg Schulenburg for providing spinpack [138], the code with which many of the computations were performed, and to Jürgen Schnack for sharing his insights on the finite-temperature Lanczos method. The $N=36$ TPQ computations were performed with a precursor of XDiag [139] using sublattice coding [140].